OpenStudy (anonymous):

Need Help with Point-Slope Form. Will Fan and Medal! Which is an equation in point-slope form for the given point and slope?

OpenStudy (anonymous):

\[Point: (2, -6) ; Slope: - \frac{3 }{ 4 }\]

OpenStudy (amistre64):

point slope .... mathmatikers arent very imaginative, they name things for the parts that they use in them. or they just blame someone else for it ..... if you cant recall the format, how do we define a slope?

OpenStudy (anonymous):

could you help me with the steps into figuring it out? I'm terrible at slopes... I always get it wrong..

OpenStudy (amistre64):

the slope formula defines the slope the slope between (2,-6) and some random general point (x,y) has to have a slope of -3/4 set up the slope formula with this information \[\frac{y--6}{x-2}=-\frac34\] we have used slope, and we have used the given point. we simply multily both sides by x-2 \[y+6=-\frac34(x-2)\] this IS the point slope format of a line.

OpenStudy (anonymous):

Thank You! :)